OpenStudy (anonymous):

Let f(x)=3x-6 and g(x)=x-2 find f/g and it's domain. My answer I got is 3: all real numbers except for x=2. Anyone care to help me?

ganeshie8 (ganeshie8):

thats \(\large \color{red}{\checkmark}\) good job !!

OpenStudy (anonymous):

I've done a few others would you mind double checking my work? Math is not my thing but i'm trying haha

ganeshie8 (ganeshie8):

you're doing great ! sure post them :)

OpenStudy (anonymous):

For the function 32/x+3 find (g^-1 times g)(5) my answer I got was 5.

ganeshie8 (ganeshie8):

\(\large \color{red}{\checkmark}\) g^-1 and g cancel each other leaving u 5

OpenStudy (anonymous):

Can I show you another? :)

ganeshie8 (ganeshie8):

shooot :)

OpenStudy (anonymous):

You can model the population of a certain city between the years 1965 and 1995 by the radical function p(x)=75,000^3 sqrt x-1940. Using this, what year would the population be 245,000. I got 1975. But, I struggled slightly so i'm not sure :/

ganeshie8 (ganeshie8):

\(\large p(x)=75,000^3 \sqrt {x-1940} \)

ganeshie8 (ganeshie8):

Is the function like above ?

OpenStudy (anonymous):

Yes :)

ganeshie8 (ganeshie8):

not sure hmm can you take a snapshot and post the exact question plzz ?

OpenStudy (anonymous):

I can't :c but I have some more questions that might be easier to answer? :)

ganeshie8 (ganeshie8):

okay !

OpenStudy (anonymous):

sqrt 2x+10^-6=2 x=27, because... s2x+10-6=2 s2x+10=2+6 2x+10=8 so square 8 take away sqrt 2x+10=64, 27 times 2 is 54 + 10 is 64

OpenStudy (anonymous):

sorry my reasoning is kind of all over the place

ganeshie8 (ganeshie8):

listen, il need a picture, its hard to interpret wat u mean by sqrt 2x+10^-6=2 :|

OpenStudy (anonymous):

Here you are c:

ganeshie8 (ganeshie8):

assuming its like below :- \(\large \sqrt{2x+10} - 6 =2 \) \(\large \sqrt{2x+10} =8\) \(\large 2x+10 =64\) \(\large 2x =54\) \(\large x =27\)

ganeshie8 (ganeshie8):

so you're rihgt !!

OpenStudy (anonymous):

I'm shocked i'm right I struggle so hard in math. This assignment though is what i'm hoping pulls my grade up :/ so you have no idea how much you're helping me out right now. Can I ask you one? I haven't found my answer yet because i'm not understand the question. I don't know if that's against this site D:

OpenStudy (anonymous):

Here's the screen shot, it's stumped me